OpenStudy (anonymous):

find f'(x) of f(x)=√(x^2 +4)

OpenStudy (jada):

f'(x)=(1/2)(x^2+4)^(-1/2)*2x=x(x^2+4)^(-1/2)

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (anonymous):

so you can use that here to get \[\frac{2x}{2\sqrt{x^2+4}}\]

OpenStudy (anonymous):

or if you like \[\frac{x}{\sqrt{x^2+4}}\]

OpenStudy (anonymous):

what if the question then ask to find the equation of the tangent line y=f(x)=√(x^2 +4) at the ponint x=1?

OpenStudy (anonymous):

well you have the point, namely 1, whatever you get when you make x=1

OpenStudy (anonymous):

i get \[(1,\sqrt{5})\]

OpenStudy (anonymous):

and you have the slope, namely whatever you get when you put x = 1 into the deivative

OpenStudy (anonymous):

make that "derivative" i get \[m=\frac{1}{\sqrt{5}}\]

OpenStudy (anonymous):

the slope is the f'(x), right?

OpenStudy (anonymous):

well \[f'(x)\] is the formula for the slope. the slope is a number. namely what you get when you replace x by whatever it is. in this case it is 1

OpenStudy (anonymous):

so the point is \[(1,\sqrt{5})\] and the slope is \[m=\frac{1}{\sqrt{5}}\] so use "point-slope" formula to get the equation for the line

OpenStudy (anonymous):

don't forget the slope is a number. the derivative is a formula for the slope, not the slope itself

OpenStudy (anonymous):

i get \[y-y_1=m(x-x_1)\] \[y-\sqrt{5}=\frac{1}{\sqrt{5}}(x-1)\] you can multiply out if you like

OpenStudy (anonymous):

ooo! Thank you! :)